OpenStudy (anonymous):

How is the following improper integral convergent? When I do partial fraction decomposition, I end up getting ln's, and if I sub in infinity into ln, I will most certainly get infinity, thus it would be divergent.... \[\int\limits_{2}^{\infty}\frac{dx}{x^2 + 2x - 3}\]

OpenStudy (anonymous):

@dpaInc

OpenStudy (rogue):

Well, looking at this, the function is smaller than 1/x^2 for x>2, so it must converge.

OpenStudy (asnaseer):

what is the partial fraction decomposition you get for this?

OpenStudy (anonymous):

Rogue is correct.

OpenStudy (asnaseer):

the integral is definitely convergent - if you show me what you did then maybe I can spot where you went wrong?

OpenStudy (rogue):

\[\int\limits_{2}^{\infty} \frac {dx}{x^2 + 2x - 3} = \frac {1}{4} \int\limits_{2}^{\infty} \frac {1}{x-1} - \frac {1}{x+3} dx = \frac {1}{4} \lim_{b \rightarrow \infty} \left[ \ln (x - 1) - \ln (x+3) \right]_{2}^{b}\]\[\frac {1}{4} \lim_{b \rightarrow \infty} \left[ \ln \frac {x - 1}{x+3} \right]_{2}^{b} = \frac {1}{4} \ln \lim_{b \rightarrow \infty} \frac {b-1}{b+3} - \frac {1}{4} \times \ln \frac {1}{4} = \frac {1}{4} ( \ln 1 - \ln \frac {1}{4}) = \frac{1}{4} ( 0 + \ln 4) = \frac {1}{4} \ln 4\]

OpenStudy (rogue):

\[= \frac {1}{4} (0 + \ln 4) = \frac {\ln 4}{4}\]

OpenStudy (asnaseer):

oh well - I guess Rogue has down all the work for you

OpenStudy (rogue):

lol, sorry :3

OpenStudy (asnaseer):

although I believe Rogue's answer is not correct

OpenStudy (rogue):

Yup, seems like I made a mistake.

OpenStudy (anonymous):

Yeah, I think there should be a 5 in there somewhere..

OpenStudy (rogue):

Yeah, I see now. It should be (ln 5) /4

OpenStudy (rogue):

When substituting 2 into x + 3, it should be 5, not 4, lol :3

OpenStudy (anonymous):

Ok I have my attempt out and ready: @asnaseer

OpenStudy (anonymous):

One esc

OpenStudy (anonymous):

I end up with those ln's at the end and those do go to infinity.

OpenStudy (anonymous):

Think you could help me with my Calc. question, Rogue?

OpenStudy (rogue):

Sure, hold on.

OpenStudy (rogue):

Apply L'hopital's rule QRAWarrior.

OpenStudy (anonymous):

How though?

OpenStudy (rogue):

Or you can just use end behavior to find the limit.

OpenStudy (asnaseer):

you can combine the two limits to get what Rogue did above

OpenStudy (anonymous):

I am going to try to combine the two limits and see what happens. Don't post any answers! I will be right back

OpenStudy (anonymous):

I think the key also is to write the integral as \[(1/4)\ln[(x-1)/(x+3)]\]

OpenStudy (asnaseer):

i.e. do this:\[\frac {1}{4} \lim_{b \rightarrow \infty} \left[ \ln (x - 1) - \ln (x+3) \right]_{2}^{b}\]

OpenStudy (asnaseer):

which in turn simplifies to:\[\frac {1}{4} \lim_{b \rightarrow \infty} \left[ \ln \frac {x - 1}{x+3} \right]_{2}^{b}\]

OpenStudy (asnaseer):

as Rogue has shown above.

OpenStudy (anonymous):

My answer is ln(5)/4.

OpenStudy (anonymous):

I took the advice of joining the logarithms together

OpenStudy (asnaseer):

perfecto!

OpenStudy (anonymous):

When I simplified the argument in the logarithm using the limit, I ended up getting ln|1|, which is 0.

OpenStudy (anonymous):

This was not even the first time I ran into the ln's and I was confused as to why it is NOT divergent

OpenStudy (anonymous):

Finally I figured it out with your guys' help

OpenStudy (anonymous):

@Rogue I wil give you a medal, if you give asnaseer a medal. I want to give you both, but I am not able to dhtat

OpenStudy (asnaseer):

I am not here for medals

OpenStudy (asnaseer):

only to help :)

OpenStudy (anonymous):

@Rogue I have one more Calc problem that involves ln's. Would you mind helping me out?

OpenStudy (rogue):

I have to go somewhere in a few minutes, sorry :( Asnaseer may be able to help :)